Number sequence:-8,-6,-4,-2 find the rule and 12th term

Answer by Edwin McCravy(20086)

-8,-6,-4,-2

The recursive rule is an = an-1 + 2, a1 = -8.

The general term is

an = a1 + (n-1)d

The common difference is 2, so d=2

an = -8 + (n-1)(2)
an = -8 + 2(n-1)
an = -8 + 2n-2
an = 2n-10
a12 = 2(12)-10 = 24-10 = 14

Checking:

-8,-6,-4,-2,0,2,4,6,8,10,12,14

Edwin